# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

### v1

@@ VI-B Map Out Hypothetical Options
Evidently, the mapping $\mathcal{F}(\cdot)$ is required to derive a variable number ($\beta$) and diverse hypothetical options for the context matrix $\{x_{i}|i\in[1,8]\}$. This indicates that conventional deep neural networks struggle to undertake the task of the mapping $\mathcal{F}(\cdot)$. Firstly, constructing a network with variable and dynamic output specifications is highly challenging. Additionally, from the perspective of neural network optimization, achieving a many-to-many mapping is infeasible. Therefore, this paper proposes a many-to-one mapping approach:
@@
Formula (23) shows that we move away from the option-mapping paradigm and instead employs a representation-mapping approach, easing the learning pressure on Brando. According to Formula (23), we have designed the structure of the Brando network as shown in the Figure 5.
@@
As indicated in Figure 5, the Brando network is formed by stacking multiple logical layers, with no parameter sharing between each layer. The input to each logical layer consists of $\{z_{ij}|i\in[1,8]\}$ and $o_{\hat{i}}$. Except for the first layer, the input to all subsequent layers comes from the output of the preceding layer. For the first layer, the $\{z_{ij}|i\in[1,8]\}$ is derived from the representations of the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution parameterized by mean and variance $N(\mu,\sigma^{2})$. For the output $o_{\hat{i}}$ from the final logic layer, this paper treats it as the representation $z_{\hat{i}j}$ of a hypothetical option provided by Brando for the context matrix $\{z_{ij}|i\in[1,8]\}$.
For the specific structure of each logical layer, it consists of a set of mapping matrices $\{W_{q},W_{k},W_{v}\}$, a multi-head self-attention mechanism, a feed-forward layer, a memory bank $M_{n}$, a buffer $B_{n}$, and a reparameterization head.
@@
    When the input set $\{z_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ is fed into a single logical layer, the mapping matrices $W_{q}$, $W_{k}$, and $W_{v}$ transform this input into query ($Q$), key ($K$), and value ($V$) representations, respectively.
@@
    After these query ($Q$), key ($K$), and value ($V$) representations are appended positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$.
@@
    As indicated by the computational step 1 in the diagram, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input representations stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the output set $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$. The learning process can be interpreted as:
@@
    In computation steps 2 and 3, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the key ($K$) and value ($V$) mappings derived from the current input representations. The update process can be expressed as:
@@
    After the learnable scalar $\tau$ is compressed into a scaling coefficient via the sigmoid function, a weighted summation is performed on $\tilde{o}_{\hat{i}}$ and $\hat{o}_{\hat{i}}$ using this coefficient to update $\hat{o}_{\hat{i}}$, enabling it to balance its own reasoning with image representations from other instances.
@@
    The attention outputs $\{\hat{z}_{ij}|i\in[1,8]\}$ with the updated $\hat{o}_{\hat{i}}$, which has undergone skip connection with logical inputs, is processed by the feed-forward layer and subjected to another skip connection, yielding the logical output $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}$. This process can be expressed as:
@@
    Here, FFL()($\cdot$) refers to the feed-forward layer.
@@
    It is noteworthy that only after the $^{o}_{\hat{i}}$ within the logical output $\{{z}_{ij},{o}_{\hat{i}}|i\in[1,8]\}$ have been processed by the reparameterization head can they serve as the final output. The process of the reparameterization head can be interpreted as follows:
@@
It is not hard to find that the memory bank $M_{n}$ and buffer $B_{n}$ in each logical layer will incorporate new memories as a new batch of context matrices is observed. Therefore, the initial form of the memory bank requires a clear definition.
@@
$M_{1}=\phi(K)^{T}V,\,B_{1}=\sum\phi(K)$ (33)
From the learnable Gaussian distribution ${N}({\mu},{\sigma}^{2})$, we sample random vectors $^{o}_{\hat{i}}$, which undergo sequential refinement in the Brando network through dynamic interaction with both current context matrix representations and historical image representations accumulated across multi-round training iterations. This iterative conditioning process enables $^{o}_{\hat{i}}$ to progressively assimilate visual patterns observed throughout the training trajectory, ultimately yielding the ultimate hypothesis representation ${z}_{\hat{i}j}$. Moreover, alterations in the initial promoter $^{o}_{\hat{i}}$ are capable of inducing changes in the ultimate option representation $z_{\hat{i}j}$. This satisfies the demand to map diverse hypothetical options for the context matrix.
@@
Prior to integrating the Brando network with DIO, DIO must be pre-trained on the original option pool $\{x_{i}|i\in[9,16]\}$ using the loss function $\ell_{\text{DIO}}$. Once DIO is pre-trained to approach its optimal performance, the Brando network will be integrated with DIO according to the method illustrated in Figure 6.
@@
Subsequently, the loss function will be switched from $\ell_{\text{DIO}}$ to $\ell_{\text{Brando}}$ for continued training of this integrated system, thereby further optimizing DIO’s performance. Another key point is that during the joint training process, DIO’s parameters excluding those in the image feature extraction module will be frozen every other iteration.
@@
It is readily apparent that the architecture of the Brando network leaves substantial room for theoretical refinement. The pursuit of a more optimal structure for implementing the mapping $\mathcal{F}(\cdot)$ appears to be an open-ended endeavor, with no discernible asymptote in sight. Moreover, providing a sufficient number of hypothetical options for DIO using Brando entails considerable computational overhead. Therefore, this paper argues that Brando is not a flawless methodology. To this end, this paper proposes a method named WORLD (Weighted Option-Represented Lower-bound Delimitation Method), which breaks away from the paradigm of the mapping $\mathcal{F}(\cdot)$. This method can also tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
VII-A Modeling is Preferable to Estimation.
Formula (12) indicates that the key to guiding the learning objective of solving model towards a tighter lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$ lies in providing additional incorrect options to the model for each context matrix. Moreover, this paper also points out that the introduced additional incorrect options must avoid being unconstructive, that is, after being incorporated into the training process of the solving model, they should be able to enhance the decision-making capability of the solving model.
@@
However, accurately estimating the distribution of RPM image features $z_{ij}$ poses significant challenges, and even if such a distribution were successfully estimated, targeted sampling from it would remain highly non-trivial. Therefore, this paper argues that explicitly modeling the distribution of $z_{ij}$ during the process where the DIO’s feature extraction module encodes $x_{i}$ into $z_{ij}$ constitutes an effective approach, since this method eliminates the need for estimation and facilitates targeted sampling through direct distributional control.
This paper has found that the Gaussian Mixture Model (GMM) [36] offers remarkable advantages, as it not only facilitates efficient modeling but also enables targeted sampling. Suppose the features $z_{ij}$ extracted by DIO from RPM images follow a GMM with known numbers of components and parameters. Then, for a given context matrix, after identifying the components corresponding to its correct options within this GMM, we can reasonably regard the components not corresponding to the correct options as the source of incorrect option features. Based on this, we are able to sample comprehensive and diverse incorrect option features from these specific components for DIO. Given these advantages, this section proposes the WORLD method to model the distribution of $z_{ij}$ as a GMM.
@@
To accurately model the distribution of RPM image features $z_{ij}$ as a GMM, this paper proposes the distribution modeling method WORLD tailored for DIO. Specifically, this paper introduces a trainable vector set $\{U_{m}|m\in[1,M]\}$ and regards it as the mean vectors of a GMM. Using this approach, we construct a predefined GMM with $M$ components. Subsequently, the image features $z_{ij}$ encoded by DIO are projected onto this predefined GMM framework, thereby modeling the distributional characteristics of $z_{ij}$.
The way in which DIO’s related modules extract features from RPM images also only requires minor modifications, so that the distribution of the extracted features $z_{ij}$ can be more easily projected onto the predefined GMM $\{U_{m}|m\in[1,M]\}$. The modifications are shown in Figure 7.
@@
The newly devised derivation method for $z_{ij}$ ensures that, when $b_{ij}$ follows a standard Gaussian distribution $\mathcal{N}(0,1)$, the distribution of $z_{ij}$ explicitly adheres to a Gaussian distribution $\mathcal{N}(u_{ij},1)$. Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the predefined GMM $\{U_{m}|m\in[1,M]\}$ and make the latent image features $b_{ij}$ follow the distribution $N(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM.
The aforementioned method for modeling the distribution of $z_{ij}$ can be implemented by introducing five additional loss function terms.
@@
    A loss term $\ell_{1}$ aims to encourage the mean of the distribution of $z_{ij}$ to approximate the predefined GMM $\{U_{m}|m\in[1,M]\}$, and its mathematical expression can be formulated as:
@@
    A loss term $\ell_{3}$, which is utilized to train the components within the predefined GMM $\{U_{m}|m\in[1,M]\}$, can be expressed as:
@@
    A mutual information loss term $\ell_{4}$, which facilitates the training of components within the predefined GMM $\{U_{m}|m\in[1,M]\}$, is expressed as:
@@
$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+0.25\cdot\ell_{3}+\ell_{4}+\ell_{5}$ (40)
By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss functions, we can train a DIO model where the distribution of the latent variable $z_{ij}$ is modeled as a GMM. Notably, when training the DIO using $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, the feedforward mechanism undergoes the transformation illustrated in Figure 8.
Figure 8: Forward Feed Process of DIO Model Incorporating $\ell_{\text{WORLD}}$ as an Additional Loss Term.
In figure 8, the latent variable fed into the progressive pattern induction module within DIO is no longer $z_{ij}$, but rather $u_{ij}+\epsilon$. In other words, the forward feed process indicated by Equation (IV-C) is updated as follows:
@@
Where $u_{ij}+\epsilon_{ij}$ can be regarded as a sample drawn from the GMM component $U_{\hat{m}}$ that $z_{ij}$ follows. This design enables DIO to interact with a fully modeled GMM rather than deterministic image features, thereby training the DIO to make decisions based on the complete probabilistic structure of the GMM instead of specific feature instantiations.
@@ VII-C Leveraging the Modeled GMM to Provide Incorrect Options for Context Matrices
where $\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2}$
$\epsilon_{m}\sim N(0,1)$
$\tau$ is a trainable smoothing parameter. Given that RPM problems often feature a diverse range of correct options, the rigid approach of designating samples drawn from components that do not correspond to the provided correct option as features of incorrect options carries a potential risk of overgeneralization. To mitigate this sample classification bias, we introduce a smoothing parameter $\tau$. $\tilde{\text{DIO}}$ refers to the DIO without the image feature extraction module, which is capable of processing features into scores. The calculation of $\tilde{m}$ serves to identify the component index in the modeled GMM $\{U_{m}|m\in[1,M]\}$ that corresponds to the feature $z_{\alpha j}$ of the correct option $x_{\alpha}$.

### v3

@@ VI-B Map Out Hypothetical Options
Evidently, the mapping $\mathcal{F}(\cdot)$ is required to derive a variable number ($\beta-17$) and diverse hypothetical options for the context matrix $\{x_{i}|i\in[1,8]\}$. This indicates that conventional deep neural networks struggle to undertake the task of the mapping $\mathcal{F}(\cdot)$. Firstly, constructing a network with variable and dynamic output specifications is highly challenging. Additionally, from the perspective of neural network optimization, achieving a many-to-many mapping is infeasible. Given these challenges, we need to explore an alternative approach. Therefore, this paper proposes a many-to-one mapping approach:
@@
Formula (23) shows that we move away from the option-mapping paradigm and instead adopt a representation-mapping approach, easing the learning pressure on Brando. According to Formula (23), we have designed the structure of the Brando network as shown in the Figure 5.
@@
As indicated in Figure 5, the Brando network is formed by stacking multiple logical layers, with no parameter sharing among each layer. The input to each logical layer consists of $\{z_{ij}|i\in[1,8]\}$ and $o_{\hat{i}}$. Except for the first layer, the input to all subsequent layers comes from the output of the preceding layer. For the first layer, $\{z_{ij}|i\in[1,8]\}$ are the features extracted by DIO from the context matrix, while the $o_{\hat{i}}$ is obtained through the reparameterization trick [2] from a learnable Gaussian distribution $N(\mu,\sigma^{2})$ parameterized by mean $\mu$ and variance $\sigma^{2}$. For the final layer, this paper treats its output $o_{\hat{i}}$ as the representation $z_{\hat{i}j}$ of a hypothetical option $x_{\hat{i}}$ provided by Brando with respect to the context matrix $\{x_{i}|i\in[1,8]\}$.
For the specific structure of each logical layer, it consists of a set of mapping matrices $\{W_{q},W_{k},W_{v}\}$, a multi-head self-attention mechanism, a feedforward layer (FFL), a dynamic memory bank $M_{n}$, a global buffer $B_{n}$, a learnable scalar $\tau$, and a reparameterization head. The feedforward process of each logical layer is as follows:
@@
    When the features $\{z_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ are input into a single logical layer, the mapping matrices $W_{q}$, $W_{k}$, and $W_{v}$ transform them into query ($Q$), key ($K$), and value ($V$) representations, respectively.
@@
    After these $Q$, $K$, and $V$ representations are appended positional encoding and processed by layer normalization, the multi-head self-attention mechanism computes interactions among them, yielding the attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$:
@@
    As indicated by the computation step 1 in the Figure 5, the normalized $Q$ representations employ a linear attention mechanism to learn the historical input feature information stored in the memory bank $M_{n-1}$ and buffer $B_{n-1}$, resulting in the memory outputs $\{\tilde{z}_{ij},\tilde{o}_{\hat{i}}|i\in[1,8]\}$. The learning process can be interpreted as:
@@
    As indicated by the computation steps 2 and 3 in the Figure 5, the memory bank $M_{n-1}$ and buffer $B_{n-1}$ are updated by leveraging the $K$ and $V$ representations derived from the current input features. The update process can be expressed as:
@@
    After the learnable scalar $\tau$ is compressed into a scaling coefficient via the sigmoid function, a weighted summation is performed on the memory output $\tilde{o}_{\hat{i}}$ and attention output $\hat{o}_{\hat{i}}$ using this coefficient. The summation result is utilized to update $\hat{o}_{\hat{i}}$, enabling it to balance current reasoning with historical feature information. The update process of $\hat{o}_{\hat{i}}$ can be expressed as:
@@
    The attention outputs $\{\hat{z}_{ij},\hat{o}_{\hat{i}}|i\in[1,8]\}$ (where $\hat{o}_{\hat{i}}$ has been updated), after undergoing residual connection with the input features, are processed by the feedforward layer and then subjected to another residual connection with their unprocessed state, ultimately yielding the logical outputs $\{z_{ij},o_{\hat{i}}|i\in[1,8]\}$. These two residual connections can be formalized as:
@@
    Importantly, the component $^{o}_{\hat{i}}$ in the logical output $\{z_{ij},o_{\hat{i}}|i\in[1,8]\}$ must be processed by the reparameterization head before serving as the final output. The reparameterization head operates as follows:
@@
As indicated in Formulas (25)-(26), in each logical layer, the memory bank $M_{n}$ and the buffer $B_{n}$ are dynamically updated as each new batch of context matrices is input. Therefore, their initial values must be explicitly defined:
@@
$M_{1}=\phi(K)^{T}V,\,B_{1}=\sum\phi(K)+10^{-9}$ (33)
From the learnable Gaussian distribution ${N}({\mu},{\sigma}^{2})$, we sample random vectors $^{o}_{\hat{i}}$, which undergo sequential refinement in the Brando network through dynamic interaction with both current context matrix representations and historical image representations accumulated over multiple rounds of training iterations. This iterative conditioning process enables $^{o}_{\hat{i}}$ to progressively assimilate visual patterns observed throughout the training trajectory, ultimately yielding the ultimate hypothesis representation ${z}_{\hat{i}j}$. Moreover, alterations in the initial promoter $^{o}_{\hat{i}}$ are capable of inducing changes in the ultimate option representation $z_{\hat{i}j}$. This satisfies the requirement to map diverse hypothetical options for the context matrix.
@@
Prior to integrating the Brando network with DIO, DIO must be pre-trained on the original option pool $\{x_{i}|i\in[9,16]\}$ using the loss function $\ell_{\text{DIO}}$. Once DIO is pre-trained to reach a state close to its optimal performance, the Brando network will be integrated with DIO according to the method illustrated in Figure 6.
@@
Subsequently, the loss function will be switched from $\ell_{\text{DIO}}$ to $\ell_{\text{Brando}}$ to continue training this integrated system, thereby further optimizing DIO’s performance. Another key point is that during the joint training process, DIO’s parameters excluding those in the image feature extraction module will be frozen every other iteration.
@@
It is readily apparent that the architecture of the Brando network leaves substantial room for theoretical refinement. The pursuit of a more optimal structure for implementing the mapping $\mathcal{F}(\cdot)$ appears to be an open-ended endeavor, with no discernible asymptote in sight. Moreover, providing a sufficient number of hypothetical options for DIO using Brando entails considerable computational overhead. Therefore, this paper argues that Brando is not a flawless method. To this end, this paper proposes a method named WORLD (Weighted Option-Represented Lower-bound Delimitation Method), which breaks away from the paradigm of the mapping $\mathcal{F}(\cdot)$. This method can also tighten the variational lower bound of mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
VII-A Local Estimation << Global Estimation << Global Modeling.
Formula (12) indicates that the key to guiding the learning objective of solving model towards a tighter lower bound of the mutual information $I(\{x_{i}\}_{i=1}^{8};x_{\alpha})$ lies in providing additional incorrect options to the model for each context matrix. Moreover, this paper also points out that the introduced additional incorrect options must avoid being unconstructive.
@@
However, accurately estimating the distribution of RPM image features $z_{ij}$ poses significant challenges, and even if such a distribution were successfully estimated, targeted sampling from it would remain highly non-trivial. Therefore, this paper argues that explicitly modeling the distribution of $z_{ij}$ during the process where the DIO’s feature extraction module encodes $x_{i}$ into $z_{ij}$ constitutes an effective approach. Because this method can obtain the distribution of $z_{ij}$ without estimation and facilitate targeted sampling by directly controlling the distribution.
This paper has found that the Gaussian Mixture Model (GMM) [36] offers remarkable advantages, as it not only facilitates efficient modeling but also enables targeted sampling. Suppose the features $z_{ij}$ extracted by DIO from RPM images follow a GMM with known numbers of components and parameters. Then, for a given context matrix, after identifying the components corresponding to its correct options within this GMM, we can reasonably regard the components not corresponding to the correct options as the source of incorrect option features. Based on this, we are able to sample comprehensive and diverse incorrect option features from these specific components for DIO to tighten its learning objective. Given these advantages, we shift our focus to transforming the underlying distribution of RPM image features into a GMM.
@@
To model the distribution of features encoded by DIO for RPM images as a GMM, this paper proposes the WORLD method. Specifically, this paper introduces a trainable vector set $\{U_{m}|m\in[1,M]\}$ and considers it as the mean vectors of a GMM. Using this approach, we define a GMM comprising $M$ components, where the means of all components are trainable while the variances remain unknown. Subsequently, by projecting the image features $z_{ij}$ encoded by DIO onto this constructed GMM framework, we can model the distributional characteristics of $z_{ij}$.
In order to enable the features $z_{ij}$ to be more easily projected onto the constructed GMM $\{U_{m}|m\in[1,M]\}$, the way DIO encodes features for RPM images requires minor modifications. The modifications are shown in Figure 7.
@@
The new derivation method for $z_{ij}$ ensures that, when $b_{ij}$ follows a standard Gaussian distribution $\mathcal{N}(0,1)$, the distribution of $z_{ij}$ explicitly adheres to a Gaussian distribution $\mathcal{N}(u_{ij},1)$. Given this, we merely need to assign the latent features $u_{ij}$ of all observed RPM images onto specific components of the constructed GMM $\{U_{m}|m\in[1,M]\}$ and make the latent features $b_{ij}$ follow the distribution $N(0,1)$, so as to effectively model the distribution of $z_{ij}$ as a GMM with trainable means and known variances.
The aforementioned method for modeling the distribution of $z_{ij}$ can be implemented by introducing five additional loss function terms:
@@
    A loss term $\ell_{1}$ aims to encourage the mean of the distribution of $z_{ij}$ to approximate the constructed GMM $\{U_{m}|m\in[1,M]\}$, and its mathematical expression can be formulated as:
@@
    A loss term $\ell_{3}$, which is utilized to train the components within the constructed GMM $\{U_{m}|m\in[1,M]\}$, can be expressed as:
@@
    A mutual information loss term $\ell_{4}$, which facilitates the training of components within the constructed GMM $\{U_{m}|m\in[1,M]\}$, is expressed as:
@@
$\ell_{\text{WORLD}}=\ell_{1}+\ell_{2}+\lambda\cdot\ell_{3}+\ell_{4}+\ell_{5}$ (40)
The coefficient $\lambda$ for $\ell_{3}$ is set to 0.25 in order to balance its relative magnitude compared to $\ell_{1}$.
By employing $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$ as the loss functions, we can train a DIO model where the distribution of image features $z_{ij}$ is modeled as a GMM. Notably, when training the DIO using $\ell_{\text{WORLD}}+\ell_{\text{DIO}}$, the feedforward mechanism undergoes the transformation illustrated in Figure 8.
Figure 8: Feedforward Process of DIO Model Incorporating $\ell_{\text{WORLD}}$ as an Additional Loss Term.
In Figure 8, the latent variable fed into the progressive pattern induction module within DIO is no longer $z_{ij}$, but rather $u_{ij}+\epsilon$. In other words, the feedforward process indicated by Equation (IV-C) is updated as follows:
@@
Where $u_{ij}+\epsilon_{ij}$ can be regarded as a sample drawn from the GMM component $U_{\hat{m}}$ that feature $z_{ij}$ follows. This design enables DIO to interact with a fully modeled GMM rather than deterministic image features, thereby training the DIO to make decisions based on the complete probabilistic structure of the GMM instead of specific feature instantiations.
@@ VII-C Leveraging the Modeled GMM to Provide Incorrect Options for Context Matrices
where $\tilde{m}=\underset{m\in[1,M]}{\text{argmin}}\|{u}_{\alpha j}-{U}_{m}\|_{2}^{2},\,\epsilon_{m}\sim N(0,1)$
$\tilde{\text{DIO}}$ refers to the DIO without the image feature extraction module, which is capable of processing image features into scores. In RPM problems, any RPM image that conforms to the progressive patterns exhibited in the context matrices can be considered a correct option. This indicates that RPM problems inherently possess a multi-solution characteristic. Consequently, the rigid approach of designating samples drawn from components that do not correspond to the provided correct option $x_{\alpha}$ as features of incorrect options carries a potential risk of overgeneralization. To mitigate this sample classification bias, we introduce a trainable smoothing parameter $\tau$. $\tilde{m}$ identifies the index of the GMM component that best matches the feature $z_{\alpha j}$ (or $u_{\alpha j}$) of the correct option $x_{\alpha}$.